OpenStudy (anonymous):

a turtle crawls at a constant speed moving 4.77 m in 19.5s . whats the velocity?

OpenStudy (ivancsc1996):

\[v=\frac{ x }{ t }=\frac{ 4.77m }{ 19.5s }\]

OpenStudy (anonymous):

a train is moving 4.82 m/s when it begins to accelerate. after 24.8s it is moving 7.09 m/s. how far did it move in that time? HELP PLZ!

OpenStudy (ivancsc1996):

\[\Delta x=v _{o}t+\frac{ 1 }{ 2 }at ^{2}=4.82\frac{ m }{ s }24.8s+\frac{ 1 }{ 2 }\frac{ 7.09\frac{ m }{ s} -4.82\frac{ m }{ s }}{ 24.8s }(24.8s)^{2}\]

OpenStudy (anonymous):

would the answer be 2.269 m/s?

OpenStudy (anonymous):

or 30.416?